when the sum of 16 and 10 times a positive number is subtracted from the sqaure of the number,o results
Answer by mananth(16946)

when the sum of 16 and 10 times a positive number is subtracted from the sqaure of the number,o results
let the number be x
x^2-(10x+16)=0
x^2-10x-16=0
find the roots of the equation
x1= (10+sqrt(100+64))/2
x1= 11.403
x2=(10-sqrt(100+64))/2
x2=-1.403
Check
plug the value of 11.403 in the equation
11.403^2-((10*11.403)+16)=.001
The number is 11.403
-1.403 does not satisfy . so ignore